### v1

Abstract
Symmetry-informed machine learning can exhibit advantages over machine learning which fails to account for symmetry. Additionally, recent attention has been given to continuous symmetry discovery using vector fields which serve as infinitesimal generators for Lie group symmetries. In this paper, we extend the notion of non-affine symmetry discovery to functions defined by neural networks. We further extend work in this area by introducing symmetry enforcement of smooth models using vector fields. Finally, we extend work on symmetry discovery using vector fields by providing both theoretical and experimental material on the restriction of the symmetry search space to infinitesimal isometries.
keywords:
Symmetry Discovery, Machine Learning, Symmetry Enforcement, Killing Vectors
{MSCcodes}
62-07, 68T01
1 Introduction
Symmetry in the context of machine learning can be exploited to improve performance and/or computational efficiency [16, 3, 7, 26]. Recently, the detection of symmetry in data has been accomplished by means of infinitesimal generators, which has allowed for the discovery of symmetries which are not affine [14, 23]. Herein, we seek to extend the work of symmetry discovery using infinitesimal generators, demonstrate a means of enforcing symmetry in machine learning models using them, and further explore the discovery of infinitesimal isometries.
Our method of discovering symmetry using infinitesimal generators is summarized as follows. First, given a dataset, certain quantities of interest, hereafter referred to as machine learning functions, are identified, such as an underlying probability distribution which generates the dataset, a classification or regression function, or a dimensionality reduction function. Second, we estimate tangent vector fields which annihilate the machine learning functions. The estimated vector fields themselves are infinitesimal generators of continuous symmetries. Thus, estimating the vector fields provides an indirect estimate of the continuous symmetries.
We offer two distinct methods of enforcing symmetry. Given previously [23], the first method uses the estimated vector fields to estimate functions which, in addition to the machine learning functions from which the vector fields were estimated, are annihilated by the vector fields. The discovered functions define a feature space which is invariant with respect to the continuous symmetries implied by the vector field generators. On this constructed feature space, the outcome of any machine learning task is, by definition, invariant with respect to the symmetries generated by the estimated vector fields. The second method of symmetry enforcement seeks to regularize the learning of a machine learning model by enforcing the vector field annihilation condition directly in the training of the model.
We also extend previous work on the restriction of the space of transformations to isometries. Previous work on symmetry discovery frequently restricts the types of symmetries being sought: common restrictions are affine symmetries [28] or classical Lie group transformations [10]. Recent work on symmetry discovery using infinitesimal generators has demonstrated potential in discovering more general transformations [14], though there is evidence that attempting to search for more generic infinitesimal generators can incur undesirable computational costs [23]. To address the shortfalls of both approaches, we propose restricting the search space to geometrically meaningful symmetries such as isometries. In principle, restrictions can be made to other types of transformations, such as conformal transformations, though at present we consider only the restriction to isometries.
This paper is organized as follows. In the next section, we discuss previous work in symmetry detection and enforcement. In section 3, we discuss vector fields, flows, and isometries, the mathematics that underlies our methods. In section 4, we describe our methods, and in section 5, we present our experimental results. We then offer concluding remarks and outline future work in the area.
2 Related Work
2.1 Symmetry Discovery
Early work on symmetry detection in machine learning focused primarily on detecting symmetry in image and video data [21, 24], where symmetries described by straight-line and rotational transformations were discovered. Other work has made strides in symmetry discovery by restricting the types of symmetries being sought. In one case, detection was limited to compact Abelian Lie groups [5] and used for the purpose of learning disentangled representations. Another case uses meta-learning to discover any finite symmetry group [30]. Finite groups have also been used in symmetry discovery in representation learning [1]. In physics-based applications of machine learning, a method has been constructed which is able to discover any classical Lie group symmetry [10]. Symmetry discovery of shapes has also been explored [12].
Other work has focused on detecting affine transformation symmetries and encoding the discovered symmetries automatically into a model architecture. Three such methods identify Lie algebra generators to describe the symmetries, as we do herein. For example, augerino [2] attempts to learn a distribution over augmentations, subsequently training a model with augmented data. The Lie algebra convolutional network [8], which generalizes Convolutional Neural Networks in the presence of affine symmetries, uses infinitesimal generators represented as vector fields to describe the symmetry. SymmetryGAN [9] has also been used to detect rotational symmetry [28].
Another notable contribution to efforts to detect symmetries of data is LieGAN. LieGAN is a generative-adversarial network intended to return infinitesimal generators of the continuous symmetry group of a given dataset [28]. LieGAN has been shown to detect continuous affine symmetries, including transformations from the Lorentz group. It has also been shown to identify discrete symmetries such as rotations by a fixed angle.
While most continuous symmetry detection methods attempt to discover symmetries which are affine transformations, the representation of infinitesimal generators using vector fields has led to the discovery of continuous symmetries which are not affine [14, 23]. In one case, the domains of image data and partial differential equations are examined in particular [14]. In the case for which this paper is an extension of [23], tabular data is examined, and the infinitesimal generators are either isometries or else usually contain polynomial coefficients.
Continuous symmetry detection is more difficult than discrete symmetry detection [30] since the condition $f\circ S=f$ must hold for all values of the continuous parameter of $S$. This is corroborated by the increasingly complex methods used to calculate even simple symmetries such as planar rotations [2, 8, 28]. Some methods introduce discretization, where multiple parameter values are chosen and evaluated. LieGAN does this by generating various transformations from the same infinitesimal generator [28]. Introducing discretization increases the complexity of continuous symmetry detection, though any reasonable symmetry detection method must establish whether a continuous symmetry a finite subgroup of a continuous group has been discovered. We believe a vector field approach addresses the issue of discretization. Our vector field-based method reduces the required model complexity of continuous symmetry detection while offering means to detect symmetries beyond affine transformations.
2.2 Symmetry Enforcement
Previous work on enforcing symmetry in machine learning models can be categorized. Some methods seek to enforce symmetry by augmenting the training dataset according to known symmetries [3]: this is common practice in image classification tasks, where augmentations include rotated copies of images in the training set. Augerino attempts to enforce symmetry using augmented data, though the symmetries are discovered from the data rather than given a priori. Another established method of enforcing symmetry is feature averaging, which is thought to be generally more effective than data augmentation [16].
A growing number of methods seek to use symmetry to construct invariant and equivariant models without the need for augmented training data, and two previously-mentioned sources do this using infinitesimal generators [8, 28]. Some previous work addresses specific cases: the special case of compact groups is studied [4], and the case of equivariant CNNs on Homegeneous spaces is studied [6]. Other work speaks to the universality of invariant architectures [17, 13, 29].
Another method which may be seen as a type of symmetry enforcement is the deployment of Physics-Informed Neural Networks (PINNs) [20]. Here, model training is regularized using differential constraints which represent the governing equations for a physical system. Our method of symmetry enforcement is, in many respects, mathematically equivalent, though the differential constrains obtained using infinitesimal generators do not generally have the interpretation of defining governing equations for a physical system.
3 Background
In this section, we will give an overview of vector fields and flows, as well as their connection to symmetry. We will also discuss Killing vectors and isometries. Throughout, we assume a dataset $\mathcal{D}=\{x_{1},\dots,x_{r}\}$ with $x_{i}\in\mathbb{R}^{n}$. Various machine learning functions of interest may exist for a given dataset, such as an underlying probability distribution or class probabilities. Generally speaking, the object of symmetry detection in data is to discover functions $S:\mathbb{R}^{n}\to\mathbb{R}^{n}$ that preserve a machine learning function of interest: that is, for a machine learning function $f$, $f\circ S=f$. We consider a continuous symmetry of a particular machine learning function to be a transformation $S$ which is continuously parameterized, such as a rotation in a plane by an angle $\theta$, with $\theta$ being the continuous parameter. We consider a discrete symmetry of a machine learning function to be a transformation which is not continuously parameterized, such as a planar rotation by a fixed angle, or a reflection of a point in the plane about a straight line. This paper deals exclusively with continuous symmetries.
3.1 Vector Fields and Flows
We now provide some background on vector fields and their associated flows (1-parameter transformations). We refer the reader to literature on the subject for additional information [15]. Suppose that $X$ is a smooth [^1] That is, a vector field with $\mathcal{C}^{\infty}$ coefficient functions. (tangent) vector field on $\mathbb{R}^{n}$:
(1) $X=\alpha^{i}\partial_{x^{i}}:=\sum_{i=1}^{n}\alpha^{i}\partial_{x^{i}},$
where $\alpha^{i}:\mathbb{R}^{n}\to\mathbb{R}$ for $i\in[1,n]$, and where $\{x^{i}\}_{i=1}^{n}$ are coordinates on $\mathbb{R}^{n}$. $X$ assigns a tangent vector at each point and can also be viewed as a function on the set of smooth, real-valued functions. E.g. if $f:\mathbb{R}^{n}\to\mathbb{R}$ is smooth,
(2) ${}X(f)=\sum_{i=1}^{n}\alpha^{i}\frac{\partial f}{\partial{x^{i}}}.$
For example, for $n=2$, if $f(x,y)=xy$ and $X=y\partial_{x}$, then $X(f)=y^{2}$. $X$ is also a derivation on the set of smooth functions on $\mathbb{R}^{n}$: that is, for smooth functions $f_{1},f_{2}:\mathbb{R}^{n}\to\mathbb{R}$ and $a_{1},a_{2}\in\mathbb{R}$,
(3) $X(a_{1}f_{1}+a_{2}f_{2})=a_{1}X(f_{1})+a_{2}X(f_{2}),\qquad X(f_{1}f_{2})=X(f_{1})f_{2}+f_{1}X(f_{2}).$
These properties are satisfied by derivatives. A flow on $\mathbb{R}^{n}$ is a smooth function $\Psi:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ which satisfies
(4) $\Psi(0,p)=p,\qquad\Psi(s,\Psi(t,p))=\Psi(s+t,p)$
for all $s,t\in\mathbb{R}$ and for all $p\in\mathbb{R}^{n}$. A flow is a 1-parameter group of transformations. An example of a flow $\Psi:\mathbb{R}\times\mathbb{R}^{2}\to\mathbb{R}^{2}$ is
(5) ${}\Psi(t,(x,y))=(x\cos(t)-y\sin(t),x\sin(t)+y\cos(t)),$
with $t$ being the continuous parameter known as the flow parameter. This flow rotates a point $(x,y)$ about the origin by $t$ radians.
For a given flow $\Psi$, one may define a (unique) vector field $X$ as given in Equation 2, where each function $\alpha^{i}$ is defined as
(6) $\alpha^{i}=\left(\frac{\partial\Psi}{\partial t}\right)\bigg{|}_{t=0}.$
Such a vector field is called the infinitesimal generator of the flow $\Psi$. For example, the infinitesimal generator of the flow given in Equation 5 is $-y\partial_{x}+x\partial_{y}$.
Conversely, given a vector field $X$ as in Equation 2, one may define a corresponding flow as follows. Consider the following system of differential equations:
(7) ${}\frac{dx^{i}}{dt}=\alpha^{i},\qquad x^{i}(0)=x^{i}_{0}.$
Suppose that a solution $\mathbf{x}(t)$ to Equation 7 exists for all $t\in\mathbb{R}$ and for all initial conditions $\mathbf{x}_{0}\in\mathbb{R}^{n}$. Then the function $\Psi:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ given by
(8) $\Psi(t,\mathbf{x}_{0})=\mathbf{x}(t)$
is a flow. The infinitesimal generator corresponding to $\Psi$ is $X$. For example, to calculate the flow of $-y\partial_{x}+x\partial_{y}$, we solve
(9) $\dot{x}=-y,\quad\dot{y}=x,\qquad x(0)=x_{0},\quad y(0)=y_{0}$
and obtain the flow $\Psi(t,(x_{0},y_{0}))$ defined by Equation 5. It is generally easier to obtain the infinitesimal generator of a flow than to obtain the flow of an infinitesimal generator.
3.2 Vector Fields and Symmetry
We can now connect vector fields and flows with symmetry. A smooth function $f:\mathbb{R}^{n}\to\mathbb{R}$ is said to be $X$-invariant if $X(f)=0$ identically for a smooth vector field $X$. The function $f$ is $\Psi$-invariant if, for all $t\in\mathbb{R}$, $f=f(\Psi(t,\cdot))$ for a flow $\Psi$. If $X$ is the infinitesimal generator of $\Psi$, $f$ is $\Psi$-invariant if and only if $f$ is $X$-invariant. If the function $f$ is a machine learning function for a given data set, our strategy is to identify vector fields $X_{i}$ for which $f$ is invariant. Each vector field $X_{i}$ for which $f$ is invariant is associated, explicitly or implicitly, with a flow $\Psi_{i}$, each of which is a 1-parameter subgroup, the collection of which generate the symmetry group. By identifying continuous symmetries by means of $X(f)=0$, continuous symmetry detection is made to be of similar complexity to that of discrete symmetry detection, since no continuous parameters are present.
The price of this indirect characterization of continuous symmetry is often the loss of the explicit identification of the symmetry group, leaving one with only the infinitesimal generators and the symmetry group merely implied. However, infinitesimal generators can be used to construct an invariant machine learning model as follows. Suppose that $f:\mathbb{R}^{n}\to\mathbb{R}$ is a (smooth) function of interest for a given data set, and suppose that $X$ is a (smooth) vector field for which $f$ is $X$-invariant. With $X$ fixed, we can consider whether there exist additional functions $\{h^{i}\}_{i=1}^{m}$ that are functionally independent of each other (meaning one function cannot be written as a function of the others), each of which are $X$-invariant. If such functions can be found, we can define a new feature space using the functions $\{h^{i}\}$. That is, we can transform each data point in $\mathcal{D}$ to the point $(h^{1}(x_{i}),h^{2}(x_{i}),\dots,h^{m}(x_{i}))$ for each point $x_{i}$ expressed in the original coordinates $\{x^{i}\}$. By definition, the transformed data is invariant with respect to the symmetry group, whether that symmetry group is given explicitly or not.
The flow parameter can often be obtained explicitly as well, written as a function of the original coordinates. It is found by solving $X(\theta)=1$. Where the flow parameters of vector fields can be estimated, our methodology thus inspires symmetry-based feature engineering, with the new features consisting of flow parameters and invariant features.
For example, consider a data set consisting of points $\mathcal{D}=\{x_{i}\}_{i=1}^{r}$ in $\mathbb{R}^{3}$, where the features of each data point are given in Cartesian coordinates $(x,y,z)$. Now suppose that each data point $x_{i}$ satisfies the equation $f=0$, where $f=x^{2}+y^{2}-z$. The function $f$ can be considered to be a machine learning function for this data set, since each point in the data set lies on the level set $f=0$. We note that $X=-y\partial_{x}+x\partial_{y}$ is a vector field for which $f$ is $X$-invariant. With $h^{1}=x^{2}+y^{2}$ and $h^{2}=z$, the functions $h^{1}$ an $h^{2}$ are $X$-invariant. We then construct the data set $\tilde{\mathcal{D}}=\{(h^{1}(x_{i}),h^{2}(x_{i})\}_{i=1}^{r}$. Any machine learning task performed in this feature space will be, by definition, invariant to the flow of $X$, which flow defines the symmetry group of the original data set $\mathcal{D}$.
Another way in which invariant machine learning models can be constructed is by imposing the condition $X(f)=0$ directly in the training of the model $f$. This is discussed in more detail in section 4.4.
By estimating vector fields which satisfy $X(f)=0$, we find seemingly unwanted freedom. This is due to the fact that, assuming $X(f)=0$, $hX(f)=0$ for a nonzero function $h$: the question, then, is which of the two symmetries–the flow of $X$ or the flow of $hX$–is “correct.” By definition, both are symmetries, so that the function $f$ is symmetric with respect to an apparently large class of symmetries. However, a function is $X$-invariant if and only if it is also invariant with respect to $hX$, so that a set of $X$-invariant functions is also a set of invariant functions of $hX$. Thus, where a vector field $X$ can be used to generate $X$-invariant functions, the result is a feature space which is also invariant with respect to $hX$. This point is further expounded upon in section 4.2.2.
3.3 Killing Vectors and Isometries
Suppose that a manifold $\mathcal{M}$ is equipped with a Riemannian metric tensor $g$. While the metric tensor is fundamentally a family of inner products defined on the tangent bundle $T\mathcal{M}$ of $\mathcal{M}$, it also induces a distance measure $d$ on $\mathcal{M}$, giving $\mathcal{M}$ the structure of a metric space. A Killing vector $X\in T\mathcal{M}$ of $g$ can be defined as a vector field which annihilates $g$ by means of the Lie derivative:
(10) ${}\mathcal{L}_{X}g=0.$
If a Killing vector $X$ is the infinitesimal generator of a flow $\Psi$, then for two points $p,q\in\mathcal{M}$,
$d(\Psi(t,p),\Psi(t,q))=d(p,q)$
for any value of the flow parameter $t$, so that $\Psi$ is an isometry of the distance measure $d$ induced by the metric tensor $g$. We note that the set of Killing vectors of a metric $g$ form a vector space as well as a Lie algebra (using the Lie bracket for vector fields). In general, Killing vectors of a particular metric $g$ can be calculated, though equation 10 defines a system of overdetermined, linear, homogeneous, partial differential equations known as the Killing equations [11], making the calculation of the Killing vectors of a metric tensor non-trivial. [^2] We note that computational methods have been developed which reduce the complexity of the Killing equations [22]. Herein, such methods are not needed, since the Killing equations for our experiments are sufficiently simple.
Within the context of machine learning, data with continuous-valued features are usually assumed not only to lie in $\mathbb{R}^{n}$, but on the Riemannian manifold $\mathbb{R}^{n}$ equipped with the Euclidean metric tensor, which metric tensor induces the Euclidean distance measure. Within this context, the Killing vectors form a proper subspace of the generators of the affine transformations. However, the Euclidean metric tensor is not always used in machine learning. This occurs when a metric tensor is learned by means of a pullback function induced from a dimensionality reduction algorithm [25]. In section 5.3, we encounter a metric tensor induced by a quadratic feature map, and we restrict our symmetry search space to vector fields which are Killing vectors.
4 Methods
Our methods assume that a given dataset consists of points residing on a differentiable manifold $\mathcal{M}$, and that a machine learning function $f$ is a smooth function on $\mathcal{M}$. We also assume that the symmetry group we seek is a Lie group generated by 1-parameter transformations known as flows, the infinitesimal generators of which are vector fields which reside in the tangent bundle of $\mathcal{M}$.
There are four components to our methodology. The first component, described in section 4.1, describes the estimation of machine learning functions specific to a dataset. The second component is the estimation of the infinitesimal symmetries of machine learning functions, as well as the estimation of invariant features of a collection of infinitesimal symmetries, and is given in section 4.2. Also in this section are two subsections which speak to the comparison of discovered symmetries to ground-truth symmetries, as well as undesired freedom present in our methods, respectively. The third component of our methodology is isometry discovery, detailed in 4.3. The final component is symmetry enforcement, discussed in section 4.4.
4.1 Estimating Machine Learning Functions
Probability density functions appear in both supervised and unsupervised problems. The output of a logistic regression model or, similarly, a neural network model that uses the softmax function, can be interpreted as estimating the probability that a given point belongs to a particular class. In fact, the probabalistic outputs of many other machine learning models can be analyzed for symmetry, provided that the probability function is differentiable. Invariant features of the symmetries of these functions define level sets upon which the probability is constant. If the flow of an infinitesimal generator can be given, a point in the dataset can be transformed to another point–not generally in the given dataset–and the model should predict the same probability for the transformed point as it would for the original point. Applications include supervised feature engineering and model validation, where predictions on generated points can be used to assess a model’s ability to generalize.
Particularly in the unsupervised case, density estimation can be challenging, complicating methods that require their estimation. A proposed method of addressing the potential shortcoming of density-based symmetry detection is called level set estimation. Level set estimation seeks to learn a function $F:\mathbb{R}^{n}\to\mathbb{R}^{k}$ ($k<n$) such that $F(x_{i})=\mathbf{0}$ for each $x_{i}\in\mathcal{D}$. Any embedded submanifold can be, at least locally, expressed as a level set of a smooth function [15], and so level set estimation operates on the assumption that the data points belonging to a given dataset lie on an embedded submanifold. If such a function can be estimated for a given data set, one can estimate symmetries of the component functions $f_{i}$ of $F$.
Level set estimation can be done as follows, although we note that further development of level set estimation may lead to more effective and/or efficient implementations. First, we construct a function $f$ as a linear combination of pre-determined features, as with a generalized additive model or polynomial regression, with model coefficients $\{a_{i}\}_{i=1}^{m}$. An $m$-component column vector $w$ is constructed using the coefficients $a_{i}$. Then, for each point in the dataset, we set the target to 00, so that a function $f$ is approximated by estimating
(11) ${}Bw=\mathcal{O},$
where $B$ is the feature matrix consisting of $m$ columns and $N$ rows, where $N$ is the number of points in the dataset, and where $\mathcal{O}$ is the zero matrix. OLS regression motivates the way in which we solve Equation (11), though we note that a solution $w_{0}$ is equivalent to a scaling $cw_{0}$ for some constant $c$. We therefore turn to constrained regression, choosing, by default, the constraint $\sum_{i=1}^{m}a_{i}^{2}=1$. Additionally, we note that any loss function can be used, with predictions and targets determined by the left- and right-hand sides of Equation (11), respectively.
We can estimate more than a single component function $f$ by appending additional columns to $w$, in which case we constrain the columns of $w$ to be orthonormal. As appending columns to $w$ is likely to increase any cost function used to estimate $w$, we use an “elbow curve” to determine when the number of columns of $w$ is sufficiently high. That is, in the presence of a sudden comparatively large increase in the cost function as the number of columns of $w$ is increased, the number of columns of $w$ is chosen as the value immediately before the large increase is encountered. We handle the constrained optimization problem of (11) using existing constrained optimization algorithms [27, 18].
The freedom of expression in the components of the level sets is, unfortunately, not fully contained by constrained optimization. When certain functions are chosen as part of the level set estimation, such as polynomials, this scaling issue can extend to multiplication by functions. For example, consider a data set in $\mathbb{R}^{3}$ known a priori to be approximately described by the function $F=\mathbf{0}$, where the components of $f_{1}$ and $f_{2}$ of $F$ are given as
(12) ${}f_{1}=x^{2}+y^{2}-1,\qquad f_{2}=z-1.$
Such a level set could also, at first glance, be described by the following equations:
(13) ${}x^{2}+y^{2}-1=0,\qquad z-1=0,\qquad x(z-1)=0.$
From a mathematical point of view, the third equation is degenerate and does not change the intrinsic dimension of the manifold characterized by this level set, owing to the fact that the rank of this function when $x=0$ is 2. However, within the context of learning level sets from data, the extrinsic dimension may be large, and it may not be practical to examine the components of the learned level set. Furthermore, key points for determining the rank of the corresponding function may not be present in the data.
A potential workaround for this problem is to incrementally increase the allowed degree of the learned polynomial functions in equation 13. For the example considered in equation 12, we can first optimize over the space of degree 1 polynomials, ideally discovering a scaling of the function $f_{2}$. Next, when searching for quadratic components, we can extend the columns of $w$ which correspond to any allowed terms of the form $h(x,y,z)f_{2}$. [^3] At degree 2, $h$ could be, for example, $x$, but not $x^{2}$, since $x^{2}z$ is not a degree 2 polynomial. With constrained optimization, the column space of $w$ can then be expressed as an orthonormal basis. The optimization routine can then continue, with the final level set model for $w$ not including the artificially-created columns. Alternatively, if one is interested primarily in at most quadratic component functions, one can first optimize using affine functions, then project the data onto the lower-dimensional hyperplane defined by the affine components, subsequently searching for quadratic terms on the reduced space where no additional affine components should be present. Experiments which apply these concepts in practice have been given previously [23].
One may well ask whether level set estimation can be accomplished using a neural network rather than constrained polynomial regression. Besides the problem of uniqueness just spoken of, a neural network approach would likely require plentiful off-manifold examples so as not to learn a function which is identically zero. A neural network approach to level set estimation is not explored in this paper, though may be explored in the future.
4.2 Estimating the Infinitesimal Generators and an Associated Invariant Feature Space
With scalar-valued machine learning functions $f_{i}$, we construct a vector-valued function $F$ whose components are defined by the functions $f_{i}$. We can then obtain vector fields which annihilate the components of $F$ by calculating nullspace vectors of $J$, the Jacobian matrix of $F$ [^4] In fact, if $F$ is given in closed form, this can be done using symbolic software and without referencing the original data. More vector fields may annihilate the functions than can be identified by the nullspace of $J$, though our method is inspired by this idea. First, we construct a vector field whose components are an arbitrary linear combination of pre-determined features, as in a general additive model, though we primarily use polynomial features herein. The coefficients of this linear combination reside in a matrix $W$, with the columns of $W$ corresponding to the coefficients for a single vector field. We then estimate $W$ in the following equation:
(14) ${}MW=\mathcal{O},$
where $M=M(\mathcal{J},B)$ is the Extended Feature Matrix computed using the array $\mathcal{J}$ of Jacobian matrices at each point, $B$ is the feature matrix, and $\mathcal{O}$ is the zero matrix. The matrix $M$ is computed via Algorithm 1.
Algorithm 1 Constructing the Extended Feature Matrix $M$
$m\leftarrow$ number of features
$n\leftarrow$ dimension of space
$N\leftarrow$ number of points in the dataset
$\mathcal{J}\leftarrow Jacobian(F)(x_{i})$triangleright 3-d array of Jacobian matrices at each point
for i in range($N$do
     $row_{0}\leftarrow padded(B_{i})$triangleright m(n-1) 0’s are appended to the $i^{th}$ row of $B$
     for j in range(1,m-1) do $row_{j}\leftarrow roll(row_{0},j\cdot m)$triangleright Each subsequent row is displaced $m$ entries to the right.
     end for
     $b_{i}\leftarrow Matrix([row_{0},\dots row_{m-1}])$
     $mat_{i}\leftarrow\mathcal{J}_{i}b_{i}$triangleright Multiply the Jacobian of $F$ at $x_{i}$ by the matrix $b_{i}$
end for
$M\leftarrow StackVertical([mat_{0},\dots mat_{N-1}])$triangleright This matrix has size $nN\times nm$.
As with Equation (11), a solution to Equation (14) is estimated using constrained optimization of a selected loss function, for which we turn to McTorch [18]. Using this pytorch-compatible library, one can choose from a variety of supported loss functions and (manifold) optimization algorithms. As with level set estimation, we sequentially increase the number of columns of $W$ until an “elbow” in cost function values is reached.
However, we find the same unwanted freedom as with level set estimation, since $X(f)=0$ implies that for a smooth function $h$, $hX(f)=0$. In the generic case, it is commonly assumed that the components of $X$ are polynomial functions, in which case it is possible that a such a pair $X$ and $hX$ may exist in the search space. The means of dealing with this unwanted freedom can depend on the particular experiment: if there are no affine symmetries, for example, vector fields with components of degree $2$ can be approximated without regard to this issue: this is further expounded upon in section 4.2.2.
To define a model which is invariant with respect to the flows of pre-computed vector fields $X_{i}$, we seek features $h^{j}$ (separate of the functions $f_{i}$) such that $X_{i}(h^{j})=0$. This we do by estimating $v$ in the following equation:
(15) ${}M_{2}v=\mathcal{O},$
where $M_{2}=M_{2}(\mathcal{J},B,\mathcal{B})$ is the Invariant Function Extended Feature Matrix computed via Algorithm 2, $\mathcal{B}$ is the feature matrix of $W$ (already estimated), and $B$ is the feature matrix of the unknown function.
Algorithm 2 Constructing the Invariant Function Extended Feature Matrix $M_{2}$
$m\leftarrow$ number of features
$n\leftarrow$ dimension of space
$N\leftarrow$ number of points in the dataset
$\bar{J}(B)(x_{i})\leftarrow$ Jacobian matrix of features of $B$ at $x_{i}$triangleright size $m\times n$: derivatives of raw features.
$\mathcal{J}\leftarrow$ $[\bar{J}(B)(x_{i})]$triangleright Array of Feature Jacobian matrices at each point.
$M=M(\mathcal{J},\mathcal{B})$triangleright Extended Feature matrix computed via Algorithm 1
$M_{2}=MW$$\triangleright$ This is not $\mathcal{O}$, since the arguments of $M$ are different than in 14
This method assumes that the invariant functions can be expressed in terms of prescribed features, which, in our case, are typically polynomial functions. The quantity $v$ can be estimated from Equation 15 in much the same way that we estimate other quantities, in that we constrain $v$ to be orthogonal, then optimize a selected loss function using McTorch [18]. We should note that while this method may produce invariant features, they are not guaranteed to be functionally independent. For example, in our polynomial feature space, the vectors corresponding to $x$ and $x^{2}$ are orthogonal, though the functions themselves are not functionally independent.
4.2.1 Comparison of Discovered Symmetries to Ground-Truth Symmetries
To evaluate symmetry discovery methods in general, it is important to quantify the ability of a given method to recover ground truth symmetries in controlled experiments. For a ground truth vector field $X$ and an estimated vector field $\hat{X}=\sum_{i=1}^{N}\hat{f}_{i}\partial_{x^{i}}$, we define the angle between them by
[…]
where $\langle X,\hat{X}\rangle_{g}=\sum_{i,j}f_{i}\hat{f}_{j}g_{ij}$, $||X||_{g}=\sqrt{\langle X,X\rangle_{g}}$, and where
[…]
with the region $\Omega$ being defined by the range of a given dataset–ordinarily, this is taken to be the full range of the dataset.
4.2.2 A Note about the Freedom Present in Identifying Vector Field Symmetries
Given a vector field $X$ and a vector field $fX$, a function $h$ is $X$-invariant if and only if it is $fX$-invariant. If $h$ is $X$-invariant, the flow of $X$ is a symmetry of $h$, so that the flow of $fX$ is also a symmetry of $h$: this follows from our discussion of vector fields in section 3.
For example, a circle centered at the origin, characterized by $F=0$ where $F(x,y)=x^{2}+y^{2}-r^{2}$ for some real number $r$, exhibits rotational symmetry described by $X=-y\partial_{x}+x\partial_{y}$. However, $\frac{1}{y}X(F)=0$, so that the flow of the vector field $-\partial_{x}+\frac{x}{y}\partial_{y}$, where defined, is also a symmetry of the circle.
This may seem to introduce a theoretical problem requiring greater care when detecting non-affine symmetry. However, our proposed methods of constructing/training models which are invariant with respect to the symmetry group requires only the identification/learning of functions which are invariant with respect to the vector fields. Thus, both $X$ and $fX$ are valid answers for “ground truth” symmetries, since a function $h$ is $X$-invariant if and only if it is $fX$-invariant.
We note that this lack of uniqueness does present a challenge when estimating suitable vector fields. When estimating the symmetries via constrained regression, $X$ and $fX$ may or may not both be present in the search space. As mentioned in section 4, one potential workaround is owing to the fact that symmetry estimation can be done symbolically. Consider the example of $F=0$ given previously in this section. The Jacobian matrix of $F$ is given as
$J=\begin{bmatrix}2x&2y&2z\end{bmatrix},$
and a basis for the nullspace of $J$ (using functions as scalars) is
$\left\{\begin{bmatrix}-y\\ x\\ 0\end{bmatrix},\begin{bmatrix}0\\ -z\\ y\end{bmatrix}\right\},$
which vectors correspond to tangent vectors $-y\partial_{x}+x\partial_{y}$ and $-z\partial_{y}+y\partial_{z}$, respectively. The flows of these vector fields correspond with rotations (about the origin) in the $(x,y)$ and $(y,z)$ planes, respectively. However, only two vector fields can be recovered using this method, which in this case has neglected another symmetry, namely $-z\partial_{x}+x\partial_{z}$, corresponding to rotations in the $(x,z)$ plane. In fact, each of the three vector fields corresponding to rotations about a coordinate axis are isometries of the sphere described by $F=0$. We should note, however, that the generator for rotations in the $(x,z)$ plane can be expressed as a linear combination of the other rotations using functions as scalars:
$\dfrac{z}{y}\left(-y\partial_{x}+x\partial_{y}\right)+\dfrac{x}{y}\left(-z\partial_{y}+y\partial_{z}\right)=-z\partial_{x}+x\partial_{z}.$
Therefore, it is plausible that applications exist in which a symbolic approach to continuous symmetry discovery may be sufficient. Our approach does not rely on symbolic software, however.
We also mention that the issue of unwanted freedom can be addressed by reducing the search space of vector fields to linear combinations of geometrically significant vector fields. In section 4.3, we discuss the restriction of the search space to Killing vectors.
Another point which may help to address concerns about our handling of the difficult nature of non-affine symmetry detection is in the relationship between the flows of $X$ and $fX$ generally. The trace of the flow of $X$ through the point $p$ is characterized by the level set $h_{i}=c_{i}$, where $\{h_{i}\}$ is a complete set of $X$-invariant functions. Since a function is $X$-invariant if and only if it is $fX$-invariant, a complete set of invariant functions for $fX$ can be taken to be $\{h_{i}\}$ without loss of generality. Thus, the level set $h_{i}=c_{i}$ also characterizes the trace of the flow of $fX$ through the point $p$. This argument assumes that the flows of $X$ and $fX$, as well as the vector fields themselves, are well-defined in an open neighborhood about the point $p$.
In fact, in our example above with $F=x^{2}+y^{2}-r^{2}$, the flow of $-\partial_{x}+\frac{x}{y}\partial_{y}$, assuming $y>0$, is given as
$\Phi(t,(x,y))=\left(x-t,\sqrt{y^{2}+2tx-t^{2}}\right),$
which trace is (part of) a circle, and we note that the flow parameter is $-x$. The trace of this flow through a point is equivalent to the trace of $X$ through the same point.
4.3 Isometry Discovery
As mentioned in section 4.2.2, the fact that a function is $X$-invariant if and only if it is also $fX$-invariant can lead to complications when estimating vector field symmetries. One way of attempting to handle this issue is to restrict the search space of symmetries so that, generally, such multiples are not present in the search space. In this section, we outline a technique to restrict the search space to isometries of a distance measure induced by a Riemannian (or pseudo-Riemannian) metric tensor.
Our model assumption is that our data lies on a manifold $\mathcal{M}$ equipped with a metric tensor $g$, and that a machine learning function $f$ on $\mathcal{M}$ is smooth. As stated before, a basis for the vector space of Killing vectors $\{X_{i}\}_{i=1}^{s}$ can be computed by solving equation 10, which is a system of overdetermined, first order, homogeneous partial differential equations known as the Killing equations. Once a basis is identified, we define $X$ to be an arbitrary linear combination of the basis vector fields:
$X=\sum_{i=1}^{s}c_{i}X_{i},$
then optimize the real coefficients $\{c_{i}\}_{i=1}^{s}$, subject to $\sum_{i=1}^{s}c_{i}^{2}=1$, in the equation $X(f)=0$. This is a case of equation 14 in which the feature matrix $B_{sr}$ is defined by the components of the Killing vectors, $\mathcal{J}$ is the Jacobian of $f$, and the single column of $W$ is defined by the coefficients $c_{i}$. To search for multiple Killing vectors which annihilate $f$, additional columns can be appended to $W$.
4.4 Symmetry Enforcement
We have discussed how, for a given smooth machine learning function $f$, one may estimate vector fields $X_{i}$ for which $X_{i}(f)=0$ identically. We now turn to the topic of enforcing continuous symmetries in a smooth machine learning model during training using vector fields.
Suppose that, in the usual way, one seeks to learn a function $f$ mapping data instances $x_{i}$ to targets $y_{i}$, where $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}^{m}$. Suppose also that the function $f$ is estimated by means of the minimization of $L(f(\mathbf{x}),\mathbf{y})$ for a smooth loss function $L$.
Now suppose that it is desired that the function $f$ learns to exhibit continuous symmetry with respect to the independent infinitesimal generators $\{X_{k}\}_{k=1}^{s}$: that is, for each component $f_{j}$ of $f$,
$X_{k}(f_{j})=0$
for $1\leq k\leq s$ and $1\leq j\leq m$. Our method of symmetry enforcement in training is to minimize the following loss function:
(17) ${}(1-\lambda(t))L(f(\mathbf{x}),\mathbf{y})+\lambda(t)\tilde{L}(\mathbf{X}(f)(\mathbf{x}),\mathcal{O}),$
where $\mathbf{X}(f)(\mathbf{x})=(X_{k}(f_{j}))(x_{i})$, $\mathcal{O}$ is an array with zero components of the same shape as $\mathbf{X}(f)(\mathbf{x})$, $\tilde{L}$ is a smooth loss function, and $\lambda(t)\in[0,1]$ is a (possibly time-dependent) symmetry regularization parameter.
The loss function given in 17 bears striking resemblance to that of a PINN, where our $\tilde{L}$ term corresponds to the appendage of differential constraints imposed by the governing equations for a given physical system [20]. Adding to this similarity is the fact that our symmetry enforcement term is a differential constraint, so that our method of symmetry enforcement appears to be a special case of a PINN in which the differential operators are linear, first-order, and homogeneous.
Despite the mathematical equivalence of enforcing continuous symmetries using vector fields and the enforcement of differential equations governing a physical system, we note that PINNs do not erase the need for symmetry enforcement using vector fields. For example, consider the enforcement of rotational symmetry in an image classification task: a symmetry-based approach provides a relatively clear path forward, since a vector field corresponding to planar rotational symmetry can readily be given. Such a task is apparently more cumbersome using a PINN, where underlying differential constraints are derived from the description of desired symmetry. Thus, a symmetry-based regularization is more appropriate when the desired symmetries can readily be given or, perhaps, more easily estimated than the full set of differential equations underlying a physical system. On the other hand, if the differential equations underlying a given system are fully known, regularization using a PINN-based approach may be the most appropriate method.
4.5 Robustness analysis
Here we present a theorem justifying the estimation of infinitesimal generators of functions that are themselves approximations of an unknown ground truth function (e.g., functions learned from data). In the theorem, $f$ plays the role of a ground truth function whose generators we wish to estimate, and $\hat{f}$ plays the role of our estimate of $f$.
Theorem 4.1 (Approximate annihilators of a function have flows that are approximate invariances).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ and $\hat{f}:\mathbb{R}^{n}\to\mathbb{R}$ be $C_{\infty}$ functions. Suppose that $\|f-\hat{f}\|_{\infty}\leq\epsilon$. Suppose that $\hat{X}$ is a vector field satisfying $\hat{X}(\hat{f})=0$ and satisfies the following Lipschitz property: for any smooth functions $g_{1},g_{2}$,
(18) $\|\hat{X}(g_{1})-\hat{X}(g_{2})\|_{\infty}\leq\lambda\cdot\|g_{1}-g_{2}\|_{\infty}.$
Let $\hat{\Psi}(t,z)$ be the flow associated with $\hat{X}$.
Then
(19) $\|\hat{X}(f)\|_{\infty}\leq\lambda\epsilon.$
As a consequence, we have, for any $t\geq 0$ and any $z\in\mathbb{R}^{n}$,
(20) $|f(\hat{\Psi}(t,z))-f(z)|\leq\lambda\epsilon t.$
Proof 4.2.
To prove (19), we write
(21) $\hat{X}(f)=\hat{X}(f-\hat{f}+\hat{f})=\hat{X}(f-\hat{f})+\hat{X}(\hat{f})=\hat{X}(f-\hat{f}),$
where the second equality is by linearity of $\hat{X}$, and the third is by the fact that $\hat{X}(\hat{f})$ annihilates $\hat{f}$.
Taking norms on both sides and using linearity of $\hat{X}$ and then the Lipschitz property,
(22) $\|\hat{X}(f)\|_{\infty}\leq\lambda\epsilon.$
To prove (20), we have
[…]
Thus, it remains to upper bound $\left|\frac{\partial f(\hat{\Psi}(\tau,z))}{\partial\tau}\right|$ for all $\tau$. We do this as follows: by the chain rule,
(24) $\frac{\partial f(\hat{\Psi}(\tau,z))}{\partial\tau}=\nabla f(\hat{\Psi}(\tau,z))\cdot\frac{\partial}{\partial\tau}\hat{\Psi}(\tau,z)=\nabla f(\hat{\Psi}(\tau,z))\cdot\hat{X}(\hat{\Psi}(\tau,z))=\hat{X}(f)(\hat{\Psi}(\tau,z)).$
The second equality is by definition of the flow associated with a vector field. The third equality is by definition of the vector field applied to functions. Applying (19), we finally get
(25) $\left|\frac{\partial f(\hat{\Psi}(\tau,z))}{\partial\tau}\right|\leq\lambda\epsilon.$
Plugging this into (23) completes the proof of the theorem.
We can interpret Theorem 4.1 as saying that a vector field that annihilates an approximation of the ground truth function $f$ has a flow that approximately serves as an invariance for $f$, where the degree of approximation depends on the smoothness of the discovered vector field and the parameter $t$ of the flow. Lipschitzness of the vector field $\hat{X}$ can be constrained by constraining it to have, e.g., polynomial coefficients with bounded degrees and monomial coefficients.
4.6 Limitations
Our methods are state of the art for continuous symmetry detection and enforcement, both for computational reasons and due to the fact that our methods can readily handle continuous symmetries which are not affine transformations and without the need to discretize continuous transformations. However, there are limitations to our methods. First, as we have mentioned previously, is the issue which arises when vector fields $X$ and $fX$ are both present in the search space. We have suggested that a symbolic approach may resolve this, though non-symbolic approaches such as ours must grapple with this issue. Our approach of searching for polynomials of lower order first partially resolves this issue. Another promising resolution is to restrict the search space of vector fields to particular types of symmetries, such as isometries, as described in section 4.3. However, this issue is not fully resolved.
Another limitation is that our method requires an assumption about the coefficient functions of the vector fields. Herein, we primarily use polynomial functions, and although polynomials are universal approximators [19], the number of polynomial terms for a given coefficient function can increase rapidly, particularly when the number of dimensions is high. We believe this limitation is best resolved, as with the previous limitation, by restricting the search space to particular types of symmetries. However, the vector fields in the restricted search space must be calculated if they are not already known.
Though our method is comparatively efficient, the reliance on the elbow curve to determine both the number of discovered symmetries and, when level set estimation is used, the number of coefficients in the level set, is a limitation. This limitation is of a more persistent nature than the previous limitations, appearing in analogous form as the dimension hyperparameter in manifold learning methods, or as the number of clusters in $k$-means clustering. It is possible that our method would be improved by another means of selecting this number, however.
We also mention the limitations inherited via the choice of optimization of equations (14) and (15). Cost function selection, optimization algorithm selection, hyperparameter selection, together with the potential for poor parameter initialization all constitute limitations to our method. We believe the optimization task is much more straight-forward than training many other symmetry detection models, however, as most other methods require fine-tuning neural networks.
Lastly, our notion of similarity of vector fields, while adequate for our experiments, is limited. Primarily, when a collection of vector fields is estimated and taken to constitute the infinitesimal generators for the symmetry group, it is not clear how to compare the estimated collection of vector fields to a ground truth collection. The similarity score also depends on the range of a given dataset, so that two vector fields may have differing similarity scores for different datasets.
5 Experiments
We offer four experiments to demonstrate our extension of previous work in the area. First, we learn the infinitesimal symmetries of a machine learning model given as a neural network: while this has been suggested as possible in principle [23], it was not explicitly accomplished. The second experiment enforces non-affine symmetry in model training, obtaining high test accuracy without the need of data augmentation. A trajectory prediction experiment follows, wherein we regularize model training using functions estimated to be invariant with respect to the Hamiltonian vector field of a given Hamiltonian. Our final experiment discovers infinitesimal isometries of a classification function, where the distance measure in the original feature space is induced by a quadratic feature map.
5.1 Symmetry Discovery for Neural Networks
In our first experiment, we demonstrate how a feedforward neural network can learn a rotationally symmetric function from 2D input data and how its learned derivatives can be used to detect the underlying symmetry. We generate $N=1000$ points $(x,y)$ uniformly from a unit disk by first sampling
$r=\sqrt{U},\quad\theta\sim\mathrm{Uniform}(0,2\pi),\quad U\sim\mathrm{Uniform}(0,1),$
and then converting to Cartesian by means of $x=r\cos(\theta)$ and $y=r\sin(\theta)$. Our regression target is $f(x,y)=x^{2}+y^{2}$, which is invariant under planar rotations about the origin. A ground-truth infinitesimal generator for this rotational symmetry is
$X=-y\partial_{x}+x\partial_{y}.$
We define a multi-layer perceptron $F:\mathbb{R}^{2}\to\mathbb{R}$ with two hidden layers of size 64 and train it for 5000 epochs using the Adam optimizer at a learning rate of 0.001. The final mean-square error (MSE) on the training set is on the order of $10^{-5}$.
To uncover the underlying rotational symmetry from the network’s derivatives, we apply equation 14 under the assumption that the coefficients of the estimated vector field $\hat{X}$ are affine. The gradient of $F$ is approximated using a finite-difference method. Using the L1 Loss, the (Riemannian) Autograd optimizer with learning rate $0.01$, and training for $5000$ epochs, we recover the following estimate of the symmetry of $F$:
$\hat{X}=(0.0012x-0.7075y+0.0002)\partial_{x}+(0.7067x-0.0058y-0.0005)\partial_{y},$
which obtains a similarity score of $0.99999$ with respect to the ground truth. This result shows that our method can estimate the infinitesimal generators of functions defined by MLPs.
5.2 Non-affine Symmetry Enforcement
Our second experiment compares ordinary polynomial regression with polynomial regression utilizing our symmetry enforcement method. First, we generate 2000 datapoints $(x_{i},y_{i})$ in the plane in a “thin strip” configuration: each $x_{i}$ is generated from $\text{Uniform}(-4,4)$, and $y_{i}=0+\epsilon_{i}$, where $\epsilon_{i}\in(-0.1,0.1)$ is a random noise component. For each point, we assign regression targets according to the function
(26) ${}f=2x^{4}-2x^{2}y^{2}+y^{4}.$
A ground truth symmetry of this function is given as
(27) ${}X=(x^{2}y-y^{3})\partial_{x}+(2x^{3}-xy^{2})\partial_{y}.$
To apply “ordinary” polynomial regression, we seek to minimize the MSE between the output of the function $f$ given in equation 26 and the output of an arbitrary degree four polynomial, iteratively updating the coefficients. We use gradient descent, a learning rate of $10^{-4}$, and $100,000$ epochs. [^5] We ignore the least squares exact solution, since this method is generally unavailable in machine learning. This results in a training MSE of $0.0159$ and a learned function given approximately as
(28) $\hat{f}=-0.0028+0.0023x-0.7872y-0.0022x^{2}+2.2454xy+0.4294y^{2}-0.0001x^{3}+0.0546x^{2}y$
$-0.5132xy^{2}+0.0426y^{3}+1.9999x^{4}-0.1950x^{3}y-0.9596x^{2}y^{2}+0.8888xy^{3}+1.7187y^{4}.$
[…]
$\cos(\theta_{\text{train}})=\dfrac{\langle\hat{f},f\rangle_{\text{train}}}{||\hat{f}||_{\text{train}}||f||_{\text{train}}}\approx 0.9999.$
[…]
$\cos(\theta_{\text{test}})=\dfrac{\langle\hat{f},f\rangle_{\text{test}}}{||\hat{f}||_{\text{test}}||f||_{\text{test}}}\approx 0.8847.$
Additionally, a visualization of the predictions for $\hat{f}$ are given in Figure 1. To estimate $f$ using symmetry enforcement, we apply equation 17 with $\lambda(t)=0.5$, with both loss functions being the MSE. Training, as before, for $100,000$ epochs, we obtain a (total) training loss function minimum of $1.4\cdot 10^{-6}$, with
(29) $\hat{f}_{s}\approx 0.0001+0.0004y^{2}+2.0000x^{4}-2.0000x^{2}y^{2}+0.9594y^{4}.$
We find that
$\cos(\theta_{\text{train}})\approx 1.0000,\qquad\cos(\theta_{\text{test}})\approx 0.9998,$
and, as Figure 1 reveals, model predictions whose symmetries appear to be more in line with that of the ground truth function.
Figure 1: Symmetry enforcement visualizations for the second experiment. Top left: an illustration of the training data and regression labels. Top right: a contour plot of the ground truth regression function. Bottom left: a contour plot of the model trained without symmetry enforcement. Bottom right: a contour plot of a model trained by enforcing symmetry.
5.3 Isometry Discovery
We consider a 3-dimensional classification problem in which points $(x,y,z)$ are assigned to class 1 if $f(x,y,z)>0$ and to class 0 otherwise, where
$f(x,y,z)=x^{2}+\tfrac{y^{2}}{2}-yz+\tfrac{z^{2}}{2}-\tfrac{1}{2}.$
In order to learn the class separation, we map the original data to a 6-dimensional Euclidean feature space only using the degree-2 terms of a polynomial feature map:
$\Phi(x,y,z)\;=\;\bigl{[}\,x^{2},\;\sqrt{2}\,x\,y,\;\sqrt{2}\,x\,z,\;y^{2},\;\sqrt{2}\,y\,z,\;z^{2}\bigr{]}.$
This mapping results in the following pullback metric on the original 3-d space:
$g=\begin{bmatrix}4x^{2}+2y^{2}+2z^{2}&2yx&2zx\\ 2yx&2x^{2}+4y^{2}+2z^{2}&2yz\\ 2zx&2yz&2x^{2}+2y^{2}+4z^{2}\end{bmatrix}.$
Next, a logistic regression classifier is trained in the 6-d feature space, producing a decision boundary that, when translated back to the original coordinates, approximately corresponds to the quadratic form $f(x,y,z)=0$. We wish to learn the isometries of the model probabilities: that is, the Killing vectors which annihilate
[…]
The metric $g$ admits three linearly independent Killing vectors, the flows of which are 3-d rotations:
$X_{1}=-\,z\,\partial_{y}+y\,\partial_{z},\quad X_{2}=-\,z\,\partial_{x}+x\,\partial_{z},\quad X_{3}=-\,y\,\partial_{x}+x\,\partial_{y}.$
We now construct an arbitrary linear combination
$\hat{X}=aX_{1}+bX_{2}+cX_{3}$
and optimize $(a,b,c)$ via equation (14). We note that a ground truth solution $X$ has coefficients $(0,-1,1)$. Using the L1 Loss function and the (Riemannian) Adagrad optimizer with a learning rate of $0.01$, we obtain the coefficients $(0.0055,0.7076,-0.7066)$, and the similarity between $X$ and $\hat{X}$ is estimated to be $0.99992$.
Thus, we have estimated the isometries of a model probability function. Because $\hat{X}$ approximately annihilates $p(x,y,z)$, the probability value of a point $\mathbf{x}$ is approximately invariant under the flow of $\hat{X}$, which is a planar rotation. Because the flow of $\hat{X}$ is an isometry, the distance (induced by the metric $g$) between two points remains unchanged when the two points are simultaneously rotated by an angle $\theta$ according to the rotation defined by the flow of $\hat{X}$.
6 Conclusion
While most current state of the art experiments focus, in the context of continuous symmetry detection, on affine symmetries, we have outlined a method to estimate symmetries including and beyond affine transformations using vector fields. Our method relies on estimating machine learning functions such as a probability distribution, continuous targets for regression, or functions which characterize an embedded submanifold (level set estimation). Vector fields are constructed as arbitrary linear combinations of a chosen basis of vector fields, and the parameters that are the coefficients of the linear combination are optimized, often using constrained optimization. Herein, we have typically expressed the coefficients of our vector fields in terms of polynomial functions.
When searching for affine symmetries in low dimensions, our method is appealing when compared with many current methods. This is due primarily to the relative ease of implementation. Additionally, when compared with the state of the art, our method appears to offer a computational advantage while at least competing and often outperforming in terms of accuracy, especially for more complex symmetries than affine ones [23].
Our infinitesimal generators are a step removed from the symmetry group than what is typical of symmetry detection methods. We have proposed building invariant architectures based on the identification of features which are invariant with respect to the discovered vector fields, resulting in an invariant model, since the information fed to the model is approximately invariant with respect to the symmetry group. We also proposed a means of symmetry enforcement in the training of a model though symmetry regularization.